# Anything Goes? A Crosslinguistic Study of (Im)possible Language Learning in LMs

### v1

@@ Abstract
Do LLMs offer insights into human language learning? A common argument against this idea is that because their architecture and training paradigm are so vastly different from humans, LLMs can learn arbitrary inputs as easily as natural languages. In this paper, we test this claim by training LMs to model impossible and typologically unattested languages. Unlike previous work, which has focused exclusively on English, we conduct experiments on 12 natural languages from 4 language families. Our results show that while GPT-2 small can primarily distinguish attested languages from their impossible counterparts, it does not achieve perfect separation between all the attested languages and all the impossible ones. We further test whether GPT-2 small distinguishes typologically attested from unattested languages with different NP orders by manipulating word order based on Greenberg’s Universal 20. We find that the model’s perplexity scores do not distinguish attested vs. unattested word orders, as long as the unattested variants maintain constituency structure. These findings suggest that language models exhibit some human-like inductive biases, though these biases are weaker than those found in human learners.
@@
Xiulin Yang1  Tatsuya Aoyama1  Yuekun Yao2  Ethan Wilcox1 1Georgetown University  2Saarland University {xy236, ta571, ethan.wilcox}@georgetown.eduykyao@coli.uni-saarland.de
@@ 1 Introduction
To what extent can LLMs serve as models of human language acquisition and processing? Some, such as Piantadosi (2023), argue that LLMs can function as comprehensive linguistic theories, challenging traditional symbolic generative approaches. On the other hand, critics maintain that the success of LLMs is largely irrelevant to human cognition due to fundamental differences in architecture and learning mechanisms (Chomsky et al., 2023; Fox and Katzir, 2024). Moreover, studies have shown that LLMs fail to acquire key aspects of linguistic knowledge, highlighting their limitations as models of human language (Fox and Katzir, 2024; Lan et al., 2024; Katzir, 2023; Dentella et al., 2024). A particularly compelling argument in this debate is that LLMs are highly flexible learners, capable of acquiring linguistic patterns beyond those learnable by humans, thus making the ability of LLMs to learn human languages uninformative for understanding human language acquisition (Chomsky and Moro, 2022; Moro, 2023; Moro et al., 2023).
@@
We present data that favors a more moderate stance, in line with other researchers in this field (Futrell and Mahowald, 2025; Millière, 2024; Pater, 2019). We present new empirical evidence from the study of impossible languages (Kallini et al., 2024) in a multilingual setting. Our findings suggest that LLMs exhibit learning biases that align with certain aspects of human cognition while simultaneously displaying biases (or a lack thereof) that diverge from human language processing.
@@
In this work, we examine both possible (attested or unattested) and impossible (unattested by definition) languages. Specifically, we define attested languages as the natural languages spoken by humans (e.g., English, German, and Chinese); unattested languages as languages constructed on language universals identified in typological studies as never-occurring; and impossible languages as those that humans cannot acquire and would never produce. Following Kallini et al. (2024), we have selected impossible variants because we take them to be uncontroversial examples of linguistic impossibility, such as languages with randomly shuffled word orders. To explore unattested languages, we draw from Greenberg’s Universal 20 (Greenberg et al., 1963), which identifies unattested word order patterns in noun phrases (e.g., adjective-number-determiner-noun). While there is no direct evidence that such languages are unlearnable, previous studies suggest that typological feature frequencies correlate with learnability in human learners (Culbertson et al., 2020; Gentner and Bowerman, 2009; Saffran et al., 2008).
@@
Regarding impossible language modeling, Kallini et al. (2024) provided initial evidence that GPT-2 small can distinguish between possible and impossible variants of English, suggesting that transformer models encode human-like linguistic biases (Futrell and Mahowald, 2025). However, their study was limited to English, leaving the question of whether this finding generalizes across languages. Furthermore, their focus on impossible languages leaves the study of unattested languages, which we take as an important testbed for LLM-human bias alignment, largely unexplored.
@@
This paper is organized around two main research questions: (1) Do LLMs encode human-like distinctions between attested and impossible languages? Specifically, (a) Within each attested language, can LLMs correctly differentiate the attested language from its impossible variants? (b) Across different attested languages from multiple language families, can LLMs distinguish all attested languages from all impossible languages? (2) Can LLMs recognize unattested languages as distinct from attested ones? Specifically, does LLMs’ ability to model unattested languages align with human typological biases?
Our findings reveal that GPT-2 small reliably distinguishes attested and impossible languages within each attested language (1a) but struggles to make this distinction across different languages (1b). It assigns lower perplexity to unattested languages when they preserve constituency and fixed word order (2), suggesting a human-like preference for regular structures (Hudson Kam and Newport, 2005; Singleton and Newport, 2004) yet deviating from the observed typological biases (e.g., Culbertson et al., 2020). [^1] Our code and data are available at https://github.com/xiulinyang/multilingual-LM.
@@ 2.1 Language Models & Cognitive Plausibility
The advancement of neural networks makes connectionism a widely adopted framework in cognitive language studies (e.g., Wilcox et al., 2023; Borenstein et al., 2024; Kirov and Cotterell, 2018). However, linguists remain divided on whether language models can meaningfully inform linguistic theories. On the one hand, language models have advanced psycholinguistics by serving as highly accurate probability estimators, and, in this capacity, have already been used for testing and refining Surprisal Theory (Goodkind and Bicknell, 2018; Oh and Schuler, 2023b, a; Kuribayashi et al., 2024), Uniform Information Density (Meister et al., 2021; Tsipidi et al., 2024), and other cognitive-linguistic theories and psychometrics (Pearl and Mis, 2011; Gibson et al., 2019; Kuribayashi et al., 2025). On the other hand, their limitations, including a lack of generalization (Yao and Koller, 2022; Kim and Linzen, 2020), the shortcomings of prompt-based approaches (Hu and Levy, 2023), and inconsistency with humans (de Dios-Flores et al., 2023; Davis and van Schijndel, 2020) suggest that, beyond their role as sophisticated estimators, they are limited as cognitive models.
The most relevant work to our study in this context is Kallini et al. (2024), which tests the hypothesis that LLMs cannot distinguish between possible and impossible languages (Chomsky et al., 2023; Moro et al., 2023). Their study relies on a 100M-word dataset from the BabyLM Challenge (Warstadt et al., 2023), focusing on systematically modified versions of English to investigate learnability and model performance. Using the language modeling task with English on GPT-2 small architecture and its impossible variants, Kallini et al. (2024) demonstrate that natural English is consistently easier to learn than its impossible counterparts, as reflected in lower perplexity scores on heldout data. They conclude that the above critique of language models as cognitive models is largely invalid.
@@ 2.2 Multilingual Language Modeling
Initial computational attempts to examine language complexity using language models were limited to RNN-based architectures (Cotterell et al., 2018; Mielke et al., 2019; Johnson et al., 2021) and $n$-grams (Koplenig and Wolfer, 2023). These studies suggest that language complexity correlates with morphological richness and the size of speaker populations. More recently, Arnett and Bergen (2025) investigated why morphologically rich languages are harder to model. By testing monolingual language models trained on carefully curated comparative datasets (Chang et al., 2024), they found that morphological features alone could not predict language learnability when training data size was controlled.
While valuable, previous studies often rely on comparative corpora, introducing inconsistencies across languages. Even with parallel corpora (Mielke et al., 2019), studies are limited by small datasets and outdated models. Our study addresses these gaps using a larger parallel corpus and modern transformer architectures.
@@ 3.1 Parallel Data Construction: OPUS12 and OPUS30
A key challenge in addressing our questions is that different languages texts drawn from different sources will have different amounts of information. To control for this, we construct two sentence-aligned multilingual parallel corpora to ensure that all languages in our dataset match in terms of content. This allows us to isolate the effect of how formal properties of a language might impact its learnability.
We name the two parallel corpora OPUS12 and OPUS30, gathering aligned sentences from five corpora available on OPUS (Tiedemann, 2012): NLLB (Schwenk et al., 2021), TED2020 (Reimers and Gurevych, 2020), the Bible (Christodouloupoulos and Steedman, 2015), OpenSubtitles (Lison and Tiedemann, 2016), and CCAligned (El-Kishky et al., 2020). Since overlap among languages decreases as more languages are included, we decide to select a minimum of 10M words in English as a standard for our parallel corpora. 10M words also correspond to the amount of input of children’s first 2 to 5 years of development (Warstadt et al., 2023).
OPUS12 is a 12-language multilingual sentence-aligned corpus [^3] The languages and their typological information are listed in Appendix C.. There are around 10M words in the case of English. OPUS30 contains 30 languages with a much smaller size: 48K sentences with 0.7M words. While the two datasets share overlapping languages, their sentences do not overlap, making OPUS30 a suitable test set for additional language modeling experiments.
@@
Table 1: Data sources of OPUS12 and OPUS30. The word counts are based on the English data. See Appendix C for licensing information.
@@ 3.3 Model Architecture & Training
In our experiments, following Kallini et al. (2024), we trained standard GPT-2 small models for each language and evaluated its performance based on the perplexity over a parallel test split of 10K randomly sampled sentences. Due to limited computational resources, we trained each model using 3 random seeds instead of the 5 used in the original study, reduced the maximum training steps from 2000 to 1200 to avoid overfitting, and adjusted the warmup steps proportionally to 120.444We did not experiment with alternative warmup steps, as Kallini et al. (2024) demonstrated that changing the warmup schedule does not affect the ranking of perplexities for impossible language models.
@@ 3.4 Multilingual Tokenization
Previous monolingual experiments either set the vocabulary size of tokenizers to be the same across languages Arnett and Bergen (2025) or applied the formula $0.4\times|V|$ Koplenig et al. (2023); Mielke et al. (2019), where $|V|$ represents the number of unique word types. We conducted a series of pilot experiments on tokenization and found neither approach suitable for our experimental design (Details can be found in Appendix B).
Given these challenges, we decided to use pretrained tokenizers. The rationale behind this choice is that when the tokenizer training data is sufficiently large and diverse, the resulting tokenization scheme should be equally good across languages, as long as the tokenizer algorithm and hyperparameters (e.g., vocabulary size, subword strategy) remain the same. [^5] Although tokenization quality, measured by metrics like compression (Schmidt et al., 2024) and Rényi entropy (Zouhar et al., 2023), has been linked to language modeling performance (e.g., Liang et al., 2023; Goldman et al., 2024), recent studies challenge this connection (Arnett and Bergen, 2025). Some may argue that the BPE algorithm might not be optimized for agglutinative languages such as Turkish, which makes the cross-linguistic comparison unfair. However, much literature on cross-linguistic LM comparison adopts BPE tokenizers (e.g., Mielke et al., 2019; Arnett and Bergen, 2025). As an additional check, we use token counts per word (TCW; reported in Appendix  E Table 5) to measure the morphological complexity of a language and report the correlation between TCW and our test-set perplexity. The results show the correlation is not significant (see Section5), suggesting that the morphological complexity of a language does not substantially impact its learnability.
While it is difficult to say how sufficiently large and diverse a tokenizer training set should be for fair comparison, we consider the size of the training data for GPT2 (Radford et al., 2019) as a reference point, as English was a high-resource language even in 2019 when the paper was published. We believe that this data size is sufficient to minimize differences that tokenization will make across languages.
When selecting pretrained tokenizers, we primarily use monolingual BPE tokenizers. [^6] However, for Chinese, we follow previous studies (Mielke et al., 2019) and use the Chinese-BERT tokenizer. Our goal is to maintain a relatively consistent vocabulary size across languages (50k), though we make exceptions for Romanian, Arabic, and Chinese due to limited model availability. The training data for all other languages is at least as large as the English corpus. The tokenizer details can be found in Appendix D.
@@
4 Experiment 1: Attested vs. Impossible Languages in an Intra-Language Modeling Setting
@@ 4.1 Impossible Languages
Since all languages are deterministically shuffled, the original ones (i.e., attested ones) can be recovered from their variants through another deterministic function. If LLMs function as non-human-like pattern recognizers as Chomsky et al. (2023); Moro et al. (2023) argue, they should be able to learn these deterministic languages.
@@ 4.2 Results & Discussion
Based on the results presented in Figure 1, we have three key observations. First, in all languages except Italian, the perplexity of the attested language is lower than all its impossible variants. For Italian, shuffle_local (w=2) yields a slightly lower perplexity than natural Italian, though the difference is not statistically significant (Mann-Whitney U test: W = 63, p = 0.353). Additionally, attested languages exhibit smaller error bars compared to their impossible counterparts, indicating more stable learning. Third, we observe a consistent learnability pattern across languages. For each language, smaller shuffling windows result in lower perplexity. Moreover, most languages perturbed with shuffle_deterministic are harder to model than shuffle_local ones. We assume this is because shuffle_deterministic languages are shuffled based on sequence length and autoregressive models do not have direct access to this information when predicting the next word, making language modeling harder.
We also conducted a Spearman’s Ranking Correlation test between the results on OPUS30 English and those from Kallini et al. (2024)’s experiments. We grouped the shuffle_deterministic languages together and observed that the ranking of our English impossible variants aligns perfectly with that reported by Kallini et al. (2024) ($\rho=1$, $p=0.0027$).
Based on these findings, we answer the first subquestion: language models can (largely) distinguish between each attested language and their corresponding impossible counterparts.
5 Experiment 2: Attested vs. Impossible Languages in an Inter-Language Modeling Setting
In this experiment, we gather the learning results of all possible and impossible languages to see if there is a separation boundary between them. If GPT-2 small can distinguish between possible and impossible languages, we expect the former’s perplexity to be lower than the latter’s.
@@
Table 3: List of NP-perturbations with corresponding categories and examples. Typo refers to typologically-attested, while Theo refers to theoretically-attested.
@@
The results of different language models are shown in Figure 2.777To highlight the overlap of perplexity between attested languages and impossible ones, we zoom in on the lower perplexity range while displaying higher perplexity values in a separate, compressed section with a break in the $y$-axis. According to the figure, we see that not every language is equally easy to learn: Chinese is the easiest language, while Arabic is the hardest, followed by Turkish and Russian. We observe a moderate positive correlation between the average number of tokens per word (TCW) and perplexity of each of the last checkpoints in 11 languages (Chinese is excluded because the BERT tokenizer is a character-level tokenizer), as indicated by a Spearman’s rank test ($\rho=0.564$), but it is not significant ($p=0.076$). This finding aligns with the observation by Arnett and Bergen (2025) that there is no significant difference in language modeling difficulty of agglutinative vs. fusional languages when the amount of information is controlled. Although all the attested languages are distributed at the bottom of the graph, we see some impossible languages fall between these attested languages. For example, Russian, Turkish and Arabic all show higher perplexity than English perturbed with shuffle_local (w=3). This means that for GPT-2 small, these impossible languages are just as hard to learn as the attested languages.
To quantify the extent GPT-2 small can distinguish attested from impossible languages, we train a linear SVM classifier with the perplexity value across the three random seeds of each checkpoint as features. The classifier reaches $0.75$ ($sd=0.08$) macro F1 score averaged over 10-folds cross-validation.
Based on this experiment, we answer the second sub-question posed in our paper: Although language models tend to learn attested languages better than impossible ones, their perplexity does not distinguish all attested languages from all impossible languages overall.
Figure 3: Attested individual natural languages vs. their corresponding impossible languages with a 95% confidence interval over 3 random seeds. The x-axis represents the training steps, and the y-axis shows the perplexity on the test split. Different language types are distinguished using distinct color palettes.
@@
In this experiment, we investigate how well language models can learn unattested languages, languages whose structure is conceivable according to rules of grammar or morphology, but which have not been attested. While unattested languages are not necessarily unlearnable (e.g., Tsimpli and Smith, 1995), prior research suggests a link between typological feature frequency, cognitive biases, and language learnability (e.g., Gentner and Bowerman, 2009; Culbertson et al., 2012; Culbertson and Newport, 2015; Culbertson et al., 2020).
We focus on Greenberg’s Universal 20 (Greenberg et al., 1963), which suggests that certain determiner-adjective-number-noun orders in an NP are universally unattested. As a well-studied typological phenomenon, Universal 20 serves as a good testbed for comparing human learners and language models, complementing experimental findings that show harmonic NP orders (i.e., ones where the dependents always all either precede or follow the head; e.g., num-adj-noun and noun-adj-num) are easier to learn than non-harmonic ones (e.g., num-noun-adj or adj-noun-num) (Culbertson and Newport, 2015, 2017; Culbertson et al., 2020). One influential hypothesis, the Typological Prevalence Hypothesis, proposes that more common typological patterns are easier to learn (Gentner and Bowerman, 2009). Therefore, we predict that if language models exhibit similar biases as humans, we expect a gradient of difficulty in learning different NP orders, with some unattested configurations posing greater challenges than others.
Among the 24 theoretically possible orders of adjectives, nouns, determiners, and numbers, we select five combinations, covering cases classified as few, many, and zero in Cinque (2005)’s typological analysis. [^8] Although Cinque (2005) seeks to explain why zero languages really are “underivable” under the minimalist program we refer to them as unattested to contrast them with the impossible languages of the previous section, i.e., ones that involve shuffling or reversed word order. In this experiment, we only permute words within NPs. If the perplexity of these permuted languages is similar to that of attested languages, it suggests two possible reasons: (1) Language models can learn these unattested languages; (2) the number of words in NP may be a small number with respect to the entire data size, and hence NP-internal perturbation introduces a much smaller noise compared to the entire data perturbation we used in previous experiments, which may not significantly affect the learnability of a language. To rule out the latter possibility, we also construct a control condition in which words corresponding to these POS categories are randomly shuffled within NPs. This language serves as a baseline, indicating the extent to which NP-internal permutations influence the learnability of a language.

### v2

@@ Abstract
Do language models (LMs) offer insights into human language learning? A common argument against this idea is that because their architecture and training paradigm are so vastly different from humans, LMs can learn arbitrary inputs as easily as natural languages. We test this claim by training LMs to model impossible and typologically unattested languages. Unlike previous work, which has focused exclusively on English, we conduct experiments on 12 languages from 4 language families with two newly constructed parallel corpora. Our results show that while GPT-2 small can largely distinguish attested languages from their impossible counterparts, it does not achieve perfect separation between all the attested languages and all the impossible ones. We further test whether GPT-2 small distinguishes typologically attested from unattested languages with different NP orders by manipulating word order based on Greenberg’s Universal 20. We find that the model’s perplexity scores do not distinguish attested vs. unattested word orders, while while its performance on the generalization does. These findings suggest that LMs exhibit some human-like inductive biases, though these biases are weaker than those found in human learners.
@@
[…]
@@ 1 Introduction
To what extent can language models (LMs) serve as models of human language acquisition and processing? Some, such as Piantadosi (2023), argue that LMs can function as comprehensive linguistic theories, challenging traditional symbolic generative approaches. However, critics maintain that the success of LMs is largely irrelevant to human cognition due to fundamental differences in architecture and learning mechanisms (Chomsky et al., 2023; Fox and Katzir, 2024). Moreover, studies have shown that LMs fail to acquire key aspects of linguistic knowledge, suggesting that they are limited as models of human language (Fox and Katzir, 2024; Lan et al., 2024; Katzir, 2023; Dentella et al., 2024). One central argument in this debate is that LMs are highly flexible learners, capable of acquiring linguistic patterns beyond those learnable by humans, thus making the ability of LMs to learn human languages uninformative for understanding human language acquisition (Chomsky and Moro, 2022; Moro, 2023; Moro et al., 2023).
@@
We present data favoring a more moderate stance, in line with other recent contributions (Futrell and Mahowald, 2025; Millière, 2024; Pater, 2019). Specifically, we present new empirical evidence from the study of impossible languages (Kallini et al., 2024) in a multilingual setting, suggesting that LMs exhibit some learning biases that align with certain aspects of human cognition. At the same time, their learning behavior is not universally human-like, suggesting that they have simultaneous biases (or a lack thereof) that diverge from human language processing.
@@
We focus on LMs’ abilities to learn different types of languages, both possible (attested or unattested) and impossible (unattested by definition). Specifically, for possible languages, we define attested languages as the natural languages spoken by humans (e.g., English, German, and Chinese); unattested languages as languages constructed on language universals and identified in typological studies as never-occurring. We consider impossible languages as those that humans cannot acquire and would never produce. Following Kallini et al. (2024), we select impossible variants as uncontroversial examples of linguistic impossibility, such as languages with shuffled or reversed word orders. To explore unattested languages, we draw from Greenberg’s Universal 20 (Greenberg et al., 1963), which identifies unattested word order patterns in noun phrases (e.g., adjective-number-determiner-noun). While there is no direct evidence that such languages are unlearnable, previous studies suggest that typological feature frequencies correlate with learnability in human learners (Culbertson et al., 2020; Gentner and Bowerman, 2009; Saffran et al., 2008).
@@
Regarding impossible language modeling, Kallini et al. (2024) provided initial evidence that GPT-2 small can distinguish between possible and impossible variants of English, suggesting that transformer models encode human-like linguistic biases (Futrell and Mahowald, 2025). However, their study was limited to English, leaving the question of whether this finding generalizes across languages unanswered. Furthermore, their focus on impossible languages leaves the study of unattested languages largely unexplored (although see Xu et al. (2025) for recent work in this area).
@@
This paper is organized around two main research questions: (1) Does LMs’ learning behavior distinguish between attested and impossible languages? Specifically, (a) Within each attested language, do LMs demonstrate better learning of an attested language compared to its impossible variants? (b) Across different attested languages from multiple language families, do LMs demonstrate better learning of all attested languages compared to all impossible languages? (2) Does LMs’ learning behavior distibguish between attested and unattested languages? Specifically, does LMs’ ability to model unattested languages align with human typological biases?
Our experiments on two parallel corpora show that GPT-2 is better at language modeling attested compared to impossible languages in most settings, though this distinction weakens for certain locally shuffled variants in some languages (1a). However, the models’ learning behavior does not distinguish attested from impossible languages across languages (1b). It assigns lower perplexity to unattested languages with preserved constituency and fixed word order, yet performs better on typologically attested languages in the generalization test (2). These findings suggest that LMs show certain human-like learning biases (e.g., Culbertson et al., 2020), though not full alignment. [^1] Our code and data are available at https://github.com/picol-georgetown/multilingual-LM.
@@ 2.1 Language Models & Cognitive Plausibility
Recent advances in deep learning have led to an upsurge in cognitive modeling with artificial neural networks, especially for language (e.g., Wilcox et al., 2023; Borenstein et al., 2024; Kirov and Cotterell, 2018). However, linguists remain divided on whether LMs can meaningfully inform linguistic theories. On the one hand, LMs are limited: They lack the capacity for (compositional) generalization (Yao and Koller, 2022; Kim and Linzen, 2020) and display biases inconsistent with human learning and processing of certain linguistic phenomena (de Dios-Flores et al., 2023; Davis and van Schijndel, 2020; Mitchell and Bowers, 2020). These issues suggest that, beyond functioning as sophisticated probability estimators, LMs have limited use as cognitive models (Cuskley et al., 2024; Bolhuis et al., 2024; Chomsky et al., 2023). Of particular relevance to our study is the argument that LMs can learn patterns that are difficult or even impossible for humans (Chomsky et al., 2023; Moro et al., 2023). This suggests that LMs do not share the cognitive constraints inherent to the human brain and may therefore miss patterns to which humans are naturally biased, rendering them uninformative for understanding human cognition.
On the other hand, LMs have advanced psycholinguistics by serving as highly accurate probability estimators and have already been used to test and refine theories such as Surprisal Theory (Goodkind and Bicknell, 2018; Oh and Schuler, 2023b, a; Kuribayashi et al., 2024), Uniform Information Density (Meister et al., 2021; Tsipidi et al., 2024), and other cognitive-linguistic theories and psychometrics (Pearl and Mis, 2011; Gibson et al., 2019; Kuribayashi et al., 2025). More recently, Kallini et al. (2024); Xu et al. (2025)’s experiments demonstrate that LMs can distinguish between possible and (typologically) impossible languages (Chomsky et al., 2023; Moro et al., 2023) in studies focusing on English and Japanese. These findings provide some empirical counter-evidence to the above arguments.
@@ 2.2 Multilingual Language Modeling
Initial computational attempts to examine language complexity using LMs were limited to RNN-based architectures (Cotterell et al., 2018; Mielke et al., 2019; Johnson et al., 2021) and $n$-grams (Koplenig and Wolfer, 2023). These studies suggest that language complexity correlates with morphological richness and the size of speaker populations. More recently, Arnett and Bergen (2025) investigated why morphologically rich languages are harder to model. By testing monolingual LMs trained on carefully curated comparative datasets (Chang et al., 2024), they found that morphological features alone could not predict language learnability when training data size was controlled.
While valuable, previous studies often rely on non-parallel corpora, introducing inconsistencies across languages. Even with parallel corpora (Mielke et al., 2019), studies are limited by small datasets and outdated models. Our study addresses these gaps using a larger parallel corpus and modern transformer architectures.
@@ 3.1 Parallel Data Construction: OPUS12 and OPUS30
One challenge in multilingual comparisons is that texts drawn from different sources in different languages will have different amounts of information. To control for this, we construct two sentence-aligned multilingual parallel corpora to ensure that all languages in our dataset match in terms of content. This allows us to isolate the effect of how formal properties of a language might impact its learnability.
We name the two parallel corpora OPUS12 and OPUS30, gathering aligned sentences from five corpora available on OPUS (Tiedemann, 2012): NLLB (Schwenk et al., 2021), TED2020 (Reimers and Gurevych, 2020), the Bible (Christodouloupoulos and Steedman, 2015), OpenSubtitles (Lison and Tiedemann, 2016), and CCAligned (El-Kishky et al., 2020). Since overlap among languages decreases as more languages are included, we decided to select a minimum of 10M words in English as a standard for our parallel corpora. 10M words also correspond to the amount of input of children’s first 2 to 5 years of development (Warstadt et al., 2023).
OPUS12 is a 12-language multilingual sentence-aligned corpus [^3] The languages and their typological information are listed in Appendix C.. There are around 10M words in the case of English. OPUS30 contains 30 languages with a smaller data size: 48K sentences with 0.7M words. While the two datasets share overlapping languages, their sentences do not overlap, making OPUS30 a suitable test set for additional language modeling experiments.
@@
Table 1: Data sources of OPUS12 and OPUS30. The word counts are based on the English data.
@@ 3.3 Model Architecture & Training
In our experiments, following Kallini et al. (2024), we trained standard GPT-2 small models for each language and evaluated its performance based on the geometric mean perplexity over a parallel test split of 10K randomly sampled sentences. Due to limited computational resources, we trained each model using 3 random seeds instead of the 5 used in the original study, reduced the maximum training steps from 2000 to 1200 to avoid overfitting, and adjusted the warmup steps proportionally to 120.444We did not experiment with alternative warmup steps, as Kallini et al. (2024) demonstrated that changing the warmup schedule does not affect the ranking of perplexities for impossible LMs.
@@ 3.4 Multilingual Tokenization
Previous monolingual experiments either set the vocabulary size of tokenizers to be the same across languages Arnett and Bergen (2025) or applied the formula $0.4\times|V|$ Koplenig et al. (2023); Mielke et al. (2019), where $|V|$ represents the number of unique word types. We conducted a series of pilot experiments on tokenization and found the latter approach unsuitable for our experimental design. Specifically, the large $|V|$ in morphologically rich languages makes it impractical to train a small model with such a large vocabulary size. Details can be found in Appendix B.
Given these considerations, we opted to use pretrained tokenizers. The rationale behind this choice is that when the tokenizer training data is sufficiently large and diverse, the resulting tokenization scheme should be equally good across languages, as long as the tokenizer algorithm and hyperparameters (e.g., vocabulary size, subword strategy) remain the same. [^5] Although tokenization quality, measured by metrics like compression (Schmidt et al., 2024) and Rényi entropy (Zouhar et al., 2023), has been linked to language modeling performance (e.g., Liang et al., 2023; Goldman et al., 2024), recent studies challenge this connection (Arnett and Bergen, 2025). While it is difficult to say how sufficiently large and diverse a tokenizer training set should be for fair comparison, we consider the size of the training data for GPT-2 (Radford et al., 2019) as a reference point, as English was a high-resource language even in 2019 when the paper was published. We believe that this data size is sufficient to minimize differences that tokenization will make across languages.
One potential concern is that the BPE algorithm might not be optimized for agglutinative languages such as Turkish. However, much literature on cross-linguistic LM comparison adopts BPE tokenizers (e.g., Mielke et al., 2019; Arnett and Bergen, 2025). As an additional check, we use token counts per word (TCW; reported in Appendix  E Table 5) to measure the morphological complexity of a language and report the correlation between TCW and our test-set perplexity. The results show the correlation is not significant (see Section5), suggesting that the morphological complexity of a language does not substantially impact its learnability in our experiments.
When selecting pretrained tokenizers, we use monolingual BPE tokenizers, [^6] However, for Chinese, we follow previous studies (Mielke et al., 2019) and use the Chinese-BERT tokenizer. targeting a vocabulary size of approximately 50k, with exceptions for Romanian, Arabic, and Chinese due to limited model availability. The training data for all other languages is at least as large as the English corpus. The tokenizer details can be found in Appendix D.
@@
4 Experiment 1: Attested vs. Impossible Languages (Intra-Language)
@@ 4.1 Impossible Languages
Since all languages are deterministically shuffled, the original ones (i.e., attested ones) can be recovered from their variants through another deterministic function. If LMs function as non-human-like pattern recognizers as Chomsky et al. (2023); Moro et al. (2023) argue, they should be able to learn these languages as well as attested ones.
@@ 4.2 Results & Discussion
The results of this experiment are presented in Figure 1. We note three high-level trends: First, in all languages except Italian, at every checkpoint, the attested language has a lower mean perplexity than all its impossible variants. For Italian, shuffle_local (w=2) yields a slightly lower perplexity than natural Italian, though the difference is not significant (Mann-Whitney U test: W = 63, p = 0.353). Welch’s t-test with Bonferroni correction across 12 checkpoints shows that for all languages, shuffle_control differs significantly from other perturbations early in training, but this difference diminishes or becomes insignificant for some languages, especially French, Italian, and Portuguese. [^7] Details in Appendix G. Attested languages also show smaller error bars, suggesting more stable learning.
Second, smaller shuffling windows consistently yield lower perplexity. Moreover, shuffle_deterministic languages result in higher perplexity than shuffle_local, likely because they shuffle based on sequence length, which autoregressive models cannot directly access. Third, as a sanity check, a Spearman’s rank correlation between OPUS30 English and Kallini et al. (2024)’s results shows strong alignment (see Appendix A).
Based on these findings, we answer the first subquestion: LMs can largely distinguish each attested language from its impossible counterparts by their learning trajectories.
5 Experiment 2: Attested vs. Impossible Languages (Inter-Language)
In this experiment, we pool the results of all possible and impossible languages and investigate whether there is a separation boundary between them. If GPT-2 small can distinguish between possible and impossible languages, we expect its perplexity on the former to be lower than on the latter.
The results of different LMs are shown in Figure 2. The first thing to note is that not every language shows the same perplexity, with Arabic highest and Chinese the lowest.
@@
Table 3: List of NP-perturbations with corresponding categories and examples. Typo refers to typologically-attested, while Theo refers to theoretically-attested by Cinque (2005)’s analysis.
@@
We observe a moderate positive correlation between the average number of tokens per word (TCW) and perplexity of each of the last checkpoints in 11 languages (Chinese is excluded because the BERT tokenizer is a character-level tokenizer), as indicated by a Spearman’s rank test ($\rho=0.564$), but it is not significant ($p=0.076$). This finding aligns with the observation by Arnett and Bergen (2025) that there is no significant difference in language modeling difficulty of agglutinative vs. fusional languages when the amount of information is controlled.
Turning to our main research question, although all the attested languages are distributed at the bottom of the graph, we see that some impossible languages fall between these attested languages. For example, Russian, Turkish, and Arabic all show higher perplexity than English perturbed with shuffle_local (w=3). To quantify the extent GPT-2’s perplexity values can separate attested from impossible languages, we train a linear SVM classifier with the perplexity value across the three random seeds of each checkpoint as features. The classifier reaches $0.75$ ($sd=0.08$) macro F1 score averaged over 10-folds cross-validation.
Based on this experiment, we answer the second sub-question posed in our paper: Although LMs tend to learn attested languages better than impossible ones, their perplexity does not distinguish all attested languages from all impossible languages.
Figure 3: Attested natural languages vs. their corresponding unattested languages with a 95% confidence interval over 3 random seeds. The x-axis represents the training steps, and the y-axis shows the perplexity on the test split. Different language types are distinguished using distinct color palettes.
@@
In this experiment, we investigate how well LMs can learn and generalize to unattested languages, languages whose structure is conceivable according to rules of grammar or morphology, but which have not been found to exist. While unattested languages are not necessarily unlearnable (e.g., Tsimpli and Smith, 1995), prior research suggests a link between typological feature frequency, cognitive biases, and language learnability (e.g., Gentner and Bowerman, 2009; Culbertson et al., 2012; Culbertson and Newport, 2015; Culbertson et al., 2020).
We focus on Greenberg’s Universal 20 (Greenberg et al., 1963), which suggests that certain determiner-adjective-number-noun orders in an NP are universally unattested. Culbertson and Newport (2015, 2017); Culbertson et al. (2020) find that harmonic NP orders (i.e., ones where the dependents always all either precede or follow the head; e.g., num-adj-noun and noun-adj-num) are easier to learn than non-harmonic ones (e.g., num-noun-adj or adj-noun-num) for humans. One influential hypothesis, the Typological Prevalence Hypothesis, proposes that more common typological patterns are easier to learn (Gentner and Bowerman, 2009). Therefore, if LMs exhibit similar biases as humans, a gradient of difficulty is expected in learning different NP orders, with some unattested configurations posing greater challenges than others.
Among the 24 theoretically possible orders of adjectives, nouns, determiners, and numbers, we select five combinations, covering cases classified as few, many, and zero in Cinque (2005)’s typological analysis. [^8] Although Cinque (2005) seeks to explain why zero languages really are “underivable” under the minimalist program we refer to them as unattested to contrast them with the impossible languages of the previous section, i.e., ones that involve shuffling or reversed word order. In this experiment, we only permute words within NPs. If the perplexity of these permuted languages is similar to that of attested languages, it suggests two possible reasons: (1) LMs can learn these unattested languages; (2) NPs may be small (in terms of number of tokens) with respect to the entire data size, and hence NP-internal perturbation introduces less noise compared to the entire data perturbation of the previous experiments. To rule out the latter possibility, we also construct a control condition in which words corresponding to these POS categories are randomly shuffled within NPs. This language serves as a baseline, indicating the extent to which NP-internal permutations influence the learnability of a language.